# Political Stress Index of Poland

### v2

@@ 1 Introduction
While Asimov’s psychohistory remains an unattainable dream, the present day large-scale, global access to vast quantities of data combined with the (so-far) exponentially increasing computational capabilities, and ever increasing sophistication of the accompanying algorithms, offer new hopes, or temptations, of forecasting human behaviour. Even if individuals remain safely unpredictable, the decisions of large social groups, could – by analogy with thermodynamics – prove amenable to mathematical analysis.
@@
Our results are threefold. First, regarding the current situation in Poland, both the original model and the one based on measured elite data do not indicate significant levels of unrest. Second, the discrepancy between the data and model led us to a deeper mathematical analysis, which shows that the equations are not well-posed, and they either have to be augmented or the interpretation of the variables has to change; we propose one such reinterpretation. Third, in an attempt to further test and calibrate the model, we applied to it to the incomplete historical data (since 1950), and the alignment with the fall of communism in Poland, despite many caveats, suggests that the model deserves further attention and refinement.
@@ 2 Political Stress Index
The components of political stress, as identified by Goldstone, are: the Mass Mobilisation Potential (MMP), Elite Mobilisation Potential (EMP) and State Fiscal Distress (SFD) (Goldstone, 1991). That they are factors contributing to unrest, revolutions or civil wars is intuitively clear, but in order to capture the information they carry, they have to be combined into a single Political Stress Index (PSI), while each component has to be either measured or further decomposed into measurable quantities.
@@ 2.1 MMP
The youth bulges numbers were taken from the Age Pyramid of GUS. Here the data also includes modelling of the population by GUS, to obtain the cohort numbers between the years when National Censuses were conducted, and up to the year 2060. For the details, the reader is directed to the official GUS methodology. Owing to the accessibility of data, we present the extended period of 1950–2023, which allows for clear identification of the “bulges”.
@@
To obtain the relative wages we take nominal values of the wages and the GDP per capita – their ratio is, by definition, unaffected by inflation, and we use nominal values in what follows in a similar manner. The resulting $w$ is shown in Figure 4.
@@ 2.2 EMP
This it the most problematic element, due to the many conflicting definition of “elite” as well as the mathematical model in the original paper (Turchin, 2013). Fortunately, we can almost completely circumvent the latter problem: there exist quite detailed data for the Polish labour market published by GUS every two years. This allows for estimation of the elite numbers and their salaries directly, rather than by modeling them based on the population-wide wages. Similar investigations have long been performed by a prominent Polish sociologist, specialising in social stratification: Henryk Domański (Domański, 2004), and we follow a similar approach.
@@ 2.3 SFD
Before jumping to the final discussion of PSI itself, we first take a historical detour, to try and reconstruct the indicators for a longer period, which lead to emergence of Poland as a truly democratic country in 1989.
@@ 3 Preliminary results: the recent years
It appears the both the data-based PSI (Figure 12) and the one obtained from the original model(Figure 13), could possibly be taken as post-revolutionary. The former is oscillatory, while the latter “exponentially” decreasing. We should be careful about this decrease, because it does not quite make sense when extended to the year 1950. That is the subject of the next section, and we will focus here on interpreting the data-based model in connection with the recent political history.
@@ 4 Excursion into the Past
Things are no quite as hopeless when it comes to GDP and debt, as there are several sources, albeit with conflicting data. This is best seen in Figure 15, which shows several reconstructions of wages relative to the GDP per capita. There is only one version of the wages data, so the various $w$ differ only through the GDP. The index had to be reconstructed, as none of the sources give nominal values in PLN for the full period. Most often international dollars are used, recalculated for specific year, with or without Purchasing Power Parity. To make them comparable with wages data, PPP had to be converted to current USD, then to current PLN, and finally, using the inflation data provided by GUS, to nominal prices. See Appendix B for sources of data, and conversion factors.
@@
Given the GDP and $w$, all other relative quantities and their derivatives follow; we show only the the elite fraction $e$, Figure 16, as it is the the central point of contention.
@@
Likewise, the national debt has been estimated by various agencies, methods and scholars. It exhibits the additional complication of the transfer rouble – an artificial currency used in obligations between countries of the communist block. The exchange rates of roubles and of Polish zlotys were of coursed fixed centrally, and diverged from the more free and in some sense more realistic black market (Kochanowski, 2009). The wild behaviour is exemplified by the Penn World Table data, which gives exchange rates in the transition period as: 0.01132 in 1984, 0.1439 in 1989, and 0.95 in 1990 – almost 2 orders of magnitude. This is further compounded with the inflation, which in 1990 reached the record level of 685.8%.
Keeping this in mind, we have pieced together the national debt as follows: for the years 1950–1970 we followed Szpringer (2012), for 1970–1980 we used Kołodko (1989), for 1980–1990 Jachowicz (2011) and IMF data, for 1990–2000 we took the raports of the Polish Ministry of Finance, and past 2000 the already mentioned GUS data. Wherever the data overlapped we took the average.
Finally, there is no reliable and continuous indicator or poll to measure public (dis)trust in the government, institutions or the state. Rather than create a patchwork approximation, we decided to exclude the variable $D$ altogether. This is not an optimal solution, but this was the scores between the past and the present become comparable.
@@ 5.1 The Dynamical Equations
If the integrand oscillates around zero, so will the integral, but that means the whole fraction, not just the wages $w$. The simplest counterexample is to take $w_{0}=1$ and $w(s)=1+a\sin(t)$ accordingly. The result, shown in Figure 19 demonstrates a catastrophic increase/decrease.
@@
will lead do increase (resp. decrease) of $e$ in time. The detailed derivation of this fact can be found in Appendix A.
@@ 5.3 The Work Ethic
The pessimism of the Polish culture, regarding an individual’s ability to overcome their plight, has been decreasing only since the end of the 20th century, when capitalistic relationships started to influence the Sarmatian ethos of Polish intelligence (Grzybowski, 1996; Cywiński, 2010). Moreover, Poland’s “youth bulges” (20–30 year-olds), react differently from those in the US, in situations perceived as unfavourable. They do not turn towards rebellion or revolution – they emigrate. Young Americans, have a much harder time finding a country to which they could escape to improve their overall situation.
@@ 5.4 Reinterpretation of $e$
We note that, the “perception” is important here, because comparison of wages in $e$ is not based on real elite income, as actual data shows. The division of surplus GDP among the elites is not the real mechanism, especially when it comes to elites that could be a conscious target for upward mobility: lawyers, doctors, or other high-level experts (like, most recently, programmers). No one can decide to change their job to “millionaire”, an old-money aristocrat, let alone backstage member of government. The job-switching decision must primarily be limited to jobs that are already included in macroeconomic statistics, and by extension in the usual relative wages $w$.
@@
In particular, for the communist Poland, the discontent is very much fueled by the inability of an individual to enter real elites or even have any democratic impact on the government. Thus without any mention of elite overproduction, the considerable growth of $e$ was still an obvious factor in the instability that would erupt in the 80’s. Especially, as pointed out in Section 5.1, we would otherwise have to accept that 93% of the late communist Poland population was in the elites (Figure 16).

### v3

@@ 1 Introduction
While Asimov’s psychohistory remains an unattainable dream, the present day large-scale, global access to vast quantities of data combined with the (so-far) exponentially increasing computational capabilities, and ever increasing sophistication of the accompanying algorithms offer new hopes, or temptations, of forecasting human behaviour. Even if individuals remain safely unpredictable, the decisions of large social groups could – by analogy with thermodynamics – prove amenable to mathematical analysis.
@@
Our results are threefold. First, regarding the current situation in Poland, both the original model and the one based on measured elite data do not indicate significant levels of unrest. Second, the discrepancy between the data and model led us to a deeper mathematical analysis, which shows that the equations are not well-posed, and they either have to be augmented or the interpretation of the variables has to change; we propose one such reinterpretation. Third, in an attempt to further test and calibrate the model, we applied it to the incomplete historical data (since 1950), and the alignment with the fall of communism in Poland, despite many caveats, suggests that the model deserves further attention and refinement.
@@ 2 Political Stress Index
The components of political stress, as identified by Goldstone, are: the Mass Mobilisation Potential (MMP), Elite Mobilisation Potential (EMP) and State Fiscal Distress (SFD) (Goldstone, 1991). That they are factors contributing to unrest, revolutions, or civil wars is intuitively clear, but in order to capture the information they carry, they have to be combined into a single Political Stress Index (PSI), while each component has to be either measured or further decomposed into measurable quantities.
@@ 2.1 MMP
The youth bulges numbers were taken from the Age Pyramid of GUS. Here, the data also include modelling of the population by GUS, to obtain the cohort numbers between the years when National Censuses were conducted, and up to the year 2060. For the details, the reader is directed to the official GUS methodology. Owing to the accessibility of data, we present the extended period of 1950–2023, which allows for clear identification of the “bulges”.
@@
To obtain the relative wages, we take nominal values of the wages and the GDP per capita – their ratio is, by definition, unaffected by inflation, and we use nominal values in what follows in a similar manner. The resulting $w$ is shown in Figure 4.
@@ 2.2 EMP
This is the most problematic element, due to the many conflicting definitions of “elite” as well as the mathematical model in the original paper (Turchin, 2013). Fortunately, we can almost completely circumvent the latter problem: there exist quite detailed data for the Polish labour market published by GUS every two years. This allows for estimation of the elite numbers and their salaries directly, rather than by modeling them based on the population-wide wages. Similar investigations have long been performed by a prominent Polish sociologist, specialising in social stratification: Henryk Domański (Domański, 2004), and we follow a similar approach.
@@ 2.3 SFD
Before jumping to the final discussion of PSI itself, we first take a historical detour, to try and reconstruct the indicators for a longer period, which lead to the emergence of Poland as a truly democratic country in 1989.
@@ 3 Preliminary results: the recent years
It appears that both the data-based PSI (Figure 12) and the one obtained from the original model(Figure 13), could possibly be taken as post-revolutionary. The former is oscillatory, while the latter “exponentially” decreasing. We should be careful about this decrease, because it does not quite make sense when extended to the year 1950. That is the subject of the next section, and we will focus here on interpreting the data-based model in connection with the recent political history.
@@ 4 Excursion into the Past
Things are not quite as hopeless when it comes to GDP and debt, as there are several sources, albeit with conflicting data. This is best seen in Figure 15, which shows several reconstructions of wages relative to the GDP per capita. There is only one version of the wages data, so the various $w$ differ only through the GDP. The index had to be reconstructed, as none of the sources give nominal values in PLN for the full period. Most often international dollars are used, recalculated for each specific year, with or without Purchasing Power Parity. To make them comparable with wages data, PPP had to be converted to current USD, then to current PLN, and finally, using the inflation data provided by GUS, to nominal prices. See Appendix B for sources of data, and conversion factors.
@@
Given the GDP and $w$, all other relative quantities and their derivatives follow; we show only the elite fraction $e$, Figure 16, as it is the central point of contention.
@@
Likewise, the national debt has been estimated by various agencies, methods and scholars. It exhibits the additional complication of the transfer rouble – an artificial currency used in obligations between countries of the communist block. The exchange rates of roubles and of Polish zlotys were of course fixed centrally, and diverged from the more free and in some sense more realistic black market (Kochanowski, 2009). The wild behaviour is exemplified by the Penn World Table data, which gives exchange rates in the transition period as: 0.01132 in 1984, 0.1439 in 1989, and 0.95 in 1990 – almost 2 orders of magnitude. This is further compounded with the inflation, which in 1990 reached the record level of 685.8%.
Keeping this in mind, we have pieced together the national debt as follows: for the years 1950–1970 we followed Szpringer (2012), for 1970–1980 we used Kołodko (1989), for 1980–1990 Jachowicz (2011) and IMF data, for 1990–2000 we took the reports of the Polish Ministry of Finance, and past 2000 the already mentioned GUS data. Wherever the data overlapped we took the average.
Finally, there is no reliable and continuous indicator or poll to measure public (dis)trust in the government, institutions or the state. Rather than create a patchwork approximation, we decided to exclude the variable $D$ altogether. This is not an optimal solution, but this way the scores between the past and the present become comparable.
@@ 5.1 The Dynamical Equations
If the integrand oscillates around zero, so will the integral, but that means the whole fraction, not just the wages $w$. The simplest counterexample is to take $w_{0}=1$ and $w(t)=1+a\sin(t)$ accordingly. The result, shown in Figure 19 demonstrates a catastrophic increase/decrease.
@@
will lead to the increase (resp. decrease) of $e$ in time. The detailed derivation of this fact can be found in Appendix A.
@@ 5.3 The Work Ethic
The pessimism of the Polish culture, regarding an individual’s ability to overcome their plight has been decreasing only since the end of the 20th century, when capitalistic relationships started to influence the Sarmatian ethos of Polish intelligence (Grzybowski, 1996; Cywiński, 2010). Moreover, Poland’s “youth bulges” (20–30 year-olds), react differently from those in the US, in situations perceived as unfavourable. They do not turn towards rebellion or revolution – they emigrate. Young Americans have a much harder time finding a country to which they could escape to improve their overall situation.
@@ 5.4 Reinterpretation of $e$
We note that the “perception” is important here, because comparison of wages in $e$ is not based on real elite income, as actual data shows. The division of surplus GDP among the elites is not the real mechanism, especially when it comes to elites that could be a conscious target for upward mobility: lawyers, doctors, or other high-level experts (like, most recently, programmers). No one can decide to change their job to “millionaire”, an old-money aristocrat, let alone a backstage member of government. The job-switching decision must primarily be limited to jobs that are already included in macroeconomic statistics, and by extension in the usual relative wages $w$.
@@
In particular, for the communist Poland, the discontent is very much fueled by the inability of an individual to enter real elites or even have any democratic impact on the government. Thus, without any mention of elite overproduction, the considerable growth of $e$ was still an obvious factor in the instability that would erupt in the 80’s. Especially, as pointed out in Section 5.1, we would otherwise have to accept that 93% of the late communist Poland population was in the elites (Figure 16).